Grade 11Mechanics

The angular speed of rotation of earth about its axis at which the weight of man standing on the equator becomes half of his weight at the poles is given by

Naman nema
7 Years agoGrade 11

1 Answer

Mohammed Kasib Siddiqui
7 Years ago
Let the weight of man at the poles = mg 
Weight at equator=mg'=1÷2 mg.
Now taking angle A=0
mg'=mg-mRw^2cosA
1÷2mg= mg-mRw^2 cos0
0.5mg= mRw^2
w^2=0.5g/R
w=(0.5g/R)^1/2 ANS.
 
Put the values of g and R to get your answer.
Plz also approve this answer.